# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ 3.3 Simulation study
To investigate the performance of the MLCM in detecting average treatment effects in panel datasets, we performed an extensive simulation study using different data-generating processes (linear and non-linear) and various combinations of pre- and post-intervention periods. In further detail, we generated $1,000$ panel datasets, each consisting of $400$ units and $T=7$ or $12$ time periods, according to the following two models,
@@
where: the error term $\epsilon_{i,t}$ is generated from a Normal distribution with standard deviation $\sigma_{\epsilon}=2$, i.e., $\epsilon_{i,t}\sim N(0,2)$, $\phi=0.8$ is the autoregressive coefficient, $\boldsymbol{\beta}=(\beta^{(1)},\dots,\beta^{(11)})^{\prime}$ is a $m\times 1$ vector of coefficients and $\mathbf{X}_{i,t-1}=\left(\operatorname{X}_{i,t-1}^{(1)},\dots,\operatorname{X}_{i,t-1}^{(11)}\right)$ is a $1\times m$ vector of $11$ predictors, both continuous and categorical, also containing interaction terms and correlated regressors. We allowed the covariates to vary in time and across units in the dataset by adding, respectively, a random term $\nu_{t}$ (which, as will become clear later, varies for different covariates) and a random term $u_{i}\sim N(1,1)$. The latter term adds variability (and thus, heterogeneity) between the units. In particular, the covariates are generated as follows: $\operatorname{X}_{i,t}^{(1)}=0.1t+u_{i}+\nu_{t}^{(1)},\nu_{t}^{(1)}\sim N(0,1)$, $\operatorname{X}_{i,t}^{(2)}=0.1t+u_{i}+\nu_{t}^{(2)},\nu_{t}^{(2)}\sim N(0,0.2)$, $\operatorname{X}_{i,t}^{(3,4,5)}=u_{i}+\nu_{t}^{(3,4,5)},\nu_{t}^{(3,4,5)}\sim MVN(0,\Sigma)$, $\operatorname{X}_{i,t}^{(6)}=u_{i}-\nu_{t}^{(6)},\nu_{t}^{(6)}\sim N(0,1)$, $\operatorname{X}_{i,t}^{(7)}=\left(0.1t+\nu_{t}^{(1)}\right)^{2}+u_{i}+\nu_{t}^{(7)},\nu_{t}^{(7)}\sim N(0,0.2)$, $\operatorname{X}_{i,t}^{(8)}\in\{0,1\}$, $\operatorname{X}_{i,t}^{(9)}\in\{1,2,3\}$, $\operatorname{X}_{i,t}^{(10)}=\operatorname{X}_{i,t}^{(3)}\cdot\operatorname{X}_{i,t}^{(9)}$, $\operatorname{X}_{i,t}^{(11)}=\operatorname{X}_{i,t}^{(2)}\cdot\operatorname{X}_{i,t}^{(8)}$. [^20] The variance-covariance matrix of the multivariate normal distribution used to generate covariates 3–5 is set to $\Sigma=\begin{bmatrix}1&.5&.7\\ .5&1&.3\\ .7&.3&1\\ \end{bmatrix}$. Notice that, to put the ML algorithms under further stress, the covariance between $\operatorname{X}_{i,t}^{(3)}$ and $\operatorname{X}_{i,t}^{(5)}$ is $0.7$ so the two variables are highly correlated but $\operatorname{X}_{i,t}^{(3)}$ is ten times more important (in terms of coefficient) than $\operatorname{X}_{i,t}^{(5)}$. These choices regarding the data-generating process for the covariates included in our simulation study are driven by the goal of mirroring real-world empirical settings where the methodology might be applied. In addition, to better reflect a typical real-world scenario where only a subset of covariates is relevant, we set certain coefficients to zero: specifically, $\beta^{(1)}$, $\beta^{(8)}$, and $\beta^{(9)}$. The remaining coefficients are generated as follows: $\beta^{(2)}=\beta^{(6)}=\beta^{(10)}=2$, $\beta^{(3)}=\beta^{(7)}=1$, $\beta^{(4)}=2.5$, $\beta^{(5)}=0.1$ and $\beta^{(11)}=1.5$. We also remark that, for computational reasons, in this simulation study we focus on a low-dimensional set of covariates. However, the MLCM is particularly well-suited for data-rich environments and sparse settings with hundreds or thousands of possible covariates. In this regard, the simulation results provided here should be interpreted as conservative evidence regarding the performance of the MLCM in real-world settings. Finally, we assume exogeneity of the predictors in both the pre- and post-intervention periods (Assumption 2). Table 1 below provides an overview of the generated datasets.
@@
At the last three time points (e.g., T = $5$, $6$, and $7$ in Table 1), we included a fictional intervention that increases the outcome for each unit by $2$ standard deviations at time $t_{0}+1$, $1.5$ standard deviations at time $t_{0}+2$, and $1$ standard deviation at time $t_{0}+3$. In other words, the intervention has a decreasing effect over time. We made this choice because adding a unit-specific component in the covariates generates heterogeneity; therefore, the scale of each $\operatorname{Y}_{i}$ varies across the $i$’s. We measured the performance of the MLCM in terms of both the bias of the estimated effect from the true impact and the interval coverage. Also note that, under this setup, the number of pre-intervention periods is $t_{0}=4$ when $T=7$ and $t_{0}=9$ when $T=12$.
The results are summarized in Table 2 below and show that, across all post-intervention horizons, the MLCM achieves a very low bias. This holds true for both linear and non-linear model specifications. The bias tends to decrease when the number of pre-intervention time periods increases, as more information is present in the data. Nevertheless, the bias at the shortest time period is still very low, which reinforces our belief that the MLCM can be effectively used for short panels. For instance, under the linear specification and $t_{0}=4$, the relative bias increases from $0.2\%$ measured at $t_{0}+1$ to $0.9\%$ measured at $t_{0}+3$. However, by increasing the number of pre-intervention time points, the bias drops at $0.4\%$ even at $t_{0}+3$. We also observe that the relative bias under a linear model specification is much lower than that under the non-linear specification, which was expected since non-linearities are typically more challenging to detect. In any case, the bias under a non-linear model specification is still low (the maximum bias is $4.9\%$ and is measured in the worst-case scenario, i.e., at the third time horizon when the pre-intervention series is very short, $t_{0}=4$).
Similar observations can be made for the interval coverage, which was estimated based on $1,000$ block-bootstrap iterations. At the first time horizon after the intervention, the interval coverage is close or equal to the nominal $95\%$ level for both linear and non-linear model specifications. Then, it tends to slightly decrease as we move further away from the intervention, but remains close to the nominal level for both linear and nonlinear processes.
@@
Notes. These results are obtained under the following scenario: $N=400$ units, $\phi=0.8$, and $\sigma_{u}=1$, where $\sigma_{u}$ is the standard deviation of the random heterogeneity term $u_{i}$. In this table, “Pre-int.” denotes the number of pre-intervention periods.
Importantly, in Appendix D, we provide many additional results that show that these key findings remain largely unchanged if we consider alternative data-generating processes for both the outcome variable and the covariates, and if we reduce the number of available units. Notably, when $\phi$ is set to $1.2$ (cf. Table 4 in Appendix D), rather than to $0.8$ as in Table 2, i.e., when moving from a stationary to a non-stationary data-generating process with an explosive autoregressive coefficient for the lag of the outcome variable, the MLCM exhibits very similar performance.
@@ 4.1 Background
In the aftermath of the COVID-19 pandemic, the inequality legacy of this unprecedented crisis has emerged as an issue of great policy relevance. The available evidence on income inequality documents a decrease driven by short-run and temporary government compensation policies, which were mostly targeted at the poorest segments of the population (Stantcheva,, 2022). Concerning education, instead, recent micro-level evidence (Agostinelli et al.,, 2022; Battisti and Maggio,, 2023; Carlana et al.,, 2023) shows that school closures had a large, persistent, and unequal effect on learning. The educational gaps caused by the school closures may, in turn, permanently affect the lifetime income possibilities of the current generation of students, with vast repercussions on future inequalities (Werner and Woessmann,, 2023). Therefore, it is through the human capital channel that the inequality effects of the pandemic may eventually appear in the medium and long run. It follows that to anticipate longer-term consequences on income inequality and territorial disparities, it is necessary to gauge the magnitude of educational losses and their distribution across areas.
To our knowledge, there is no granular evidence on the geography of the education effects of the COVID-19 crisis. This is mainly due to econometric challenges caused by the sudden spread of the pandemic across the world, which resulted in the absence of a suitable untreated group. Italy is an important case study, as it ranks among the hardest-hit countries and was the first Western country to impose a strict nationwide lockdown. During the lockdown that began on March 9, 2020, schools across the entire national territory were closed and remained so until the end of the school year. This resulted in Italy ranking among the OECD countries with the highest number of weeks of school closures and distance learning (Battisti and Maggio,, 2023). In this setting, the treatment—the COVID-19 shock—simultaneously affected all units. [^21] We refer to the treatment variable as COVID-19 “shock”, which encompasses both the pandemic and the containment policies, as we aim to capture the total effect of the pandemic, i.e., its direct and indirect effects on educational outcomes. In this regard, in line with previous literature on COVID-19, we consider the restrictive measures, including the lockdown and the school closures, as a manifestation of the pandemic, not as a separate treatment. COVID-related learning losses can originate from many pandemic-induced channels, such as school closures and distance learning, prolonged absence from school due to one or more infections, absence due to parents’ concerns about their children’s health, and many other transmission mechanisms. At the aggregate level, only the overall impact of the pandemic remains discernible. Therefore, we leverage the MLCM.
@@ 4.2 Data and implementation
We employ LLM yearly data covering the period from 2013 to 2020. We cover all Italian LLMs except some of the smallest ones for which the education data are unavailable due to privacy protection, for a total of $579$ LLMs ($95\%$ of Italian LLMs and over $99\%$ of the Italian population). The dependent variable is the standardized math test score of fifth-grade students referred to the school year started in 2020. Following the approach by Carlana et al., (2023), the math score is standardized with respect to its pre-pandemic mean. This way, the detected treatment effects can be interpreted as deviations from the pre-COVID baseline. We focus on younger students for two reasons: i) learning disruptions earlier in life typically have longer-lasting and more severe effects, so younger children may have been more heavily impacted (Stantcheva,, 2022); ii) primary schools were excluded from the heterogeneous school closures involved in the tier system of regional restrictions implemented by the Italian government since the onset of the second wave (November 2020), so the treatment is the same for all units (Assumption 1). The initial pre-treatment information set includes over $150$ variables (see Table 11 in Appendix E for a detailed description of the variables). In this set of covariates, we included the first three lags of all the predictors as covariates and three lags of the outcome variable. This implies that we collapse the original 2013–2020 dataset into a dataset covering the period 2017–2020.
@@
We then use a data-driven approach to restrict the information set. More specifically, we follow Athey and Wager, (2019) and Basu et al., (2018) and apply a pilot random forest on the pre-treatment data to pick up a subset of the most important predictors according to the importance ranking produced by the forest. [^22] For this preliminary operation, we use default hyperparameter settings that typically perform well with random forests (Athey and Wager,, 2019). In order to select the precise number of relevant predictors in a data-driven manner, we include this number as an additional parameter in the subsequent panel CV routine which we employ to tune the hyperparameters of all the selected algorithms (LASSO, Partial Least Squares, stochastic gradient boosting, and random forest, see Appendix B for an overview of these models). [^23] For all algorithms, we tune the subset of most predictive variables (10, 20 or 30) to be included in the analysis according to the preliminary forest. For LASSO, we tune the penalty parameter (the values from 0.1 to 0.9 with an increase of 0.1 in each run); for PLS, we tune the number of components (all possible values from 1 to the number of variables). For boosting, we tune the number of trees (1,000 or 2,000), the maximum depth of each tree (1 or 2), the minimum number of observations in terminal nodes (5 or 10) and the learning rate (0.001, 0.002 or 0.005); for random forest, we tune the number of variables randomly sampled as candidates at each split (1/2, 1/3 or 1/4 of the number of variables), and use a fixed number of 1,000 trees. All candidate hyperparameter values are tested on each testing set. The implementation process for this application is reported in Box 4.2 below, while the outcomes of the empirical analysis are reported in Section 4.3. Preliminary   0. Data splitting. We split the full 2017-2020 dataset into two subsets according to the treatment date: 2017-2019 to be used in the Design Stage; 2020 to be used in the Analysis Stage.   Design Stage   1. Algorithm selection. We select four supervised ML algorithms: 1) LASSO; 2) Partial Least Squares; 3) stochastic gradient boosting; 4) random forest. As we are agnostic about the functional form of the underlying data-generating process, we opt for a mix of non-linear and linear models. 2. Principled input selection. We build an initial LLM dataset with over $150$ predictors on the basis of literature insights. From this dataset, we then keep only the most important predictors selected by a preliminary random forest run on the pre-treatment data (see step below). 3. Panel cross-validation. For each algorithm, we tune hyperparameters via panel CV, involving iterative estimation on two different training-testing pairs of pre-COVID datasets: i) 2017 training; 2018 testing; ii) 2017-2018 training, 2019 testing. 4. Performance assessment. We assess average performance metrics for the four algorithms by comparing average forecasted vs. actual outcomes on the 2018–2019 held-out test data. We then compare the performance of the different MLCM versions. 5. Diagnostic and placebo tests. We first check the average distribution of errors with the best-performing model for the 2018-2019 testing sets and then show the map of the unit-level placebo temporal average treatment effects in the pre-COVID period.   Analysis Stage   6. Final model selection. On the basis of the comparative performance assessment, we pick the best-performing algorithm (random forest) and re-train it on the full 2017–2019 sample using the hyperparameters cross-validated in the Design stage. 7. Counterfactual forecasting. We apply the final model estimated in Step 6 on the post-pandemic data and forecast, for each LLM $i$, the counterfactual outcome $\hat{\operatorname{Y}}_{i,t_{0}+1}$. 8. Estimation of treatment effects. For each LLM $i$, we estimate the individual treatment effect by taking the difference between the observed post-COVID outcome $\operatorname{Y}_{i,t_{0}+1}$ and the ML-generated potential outcome $\hat{\operatorname{Y}}_{i,t_{0}+1}$. 9. Treatment effect heterogeneity. We estimate data-driven CATEs via a regression tree analysis with the individual treatment effects as the outcome and a host of potentially relevant predictors associated with the heterogeneity of the education effects. 10. Inference. We estimate standard errors for the ATE and CATEs via block-bootstrapping by performing $1,000$ bootstrap replications of Steps 6 to 9.
@@
Note that in this application, due to the ubiquitous nature of this shock, we do not invoke the second part of Assumption 2 (exogeneity of post-treatment covariates) and only include pre-treatment values of time-varying variables and time-invariant predictors in the information set. However, given the unprecedented disruptions brought about by the pandemic and based on institutional and subject matter knowledge, we deem it plausible that any divergence from the forecasted counterfactual in educational outcomes can be attributed exclusively to the COVID-19 shock, and not to other concomitant shocks or policies. [^24] There were no educational reforms at the primary school level that could have influenced the standardized math test scores of fifth-grade students in the period under scrutiny.

### v2

@@ 3.3 Simulation study
To investigate the performance of the MLCM in detecting average treatment effects in panel datasets, we performed an extensive simulation study using different data-generating processes (linear and non-linear) and various combinations of pre- and post-intervention periods. We generated $1,000$ panel datasets, each consisting of $400$ units and $T=7$ or $12$ time periods, according to the following two models,
@@
where: the error term $\epsilon_{i,t}$ is generated from a Normal distribution with standard deviation $\sigma_{\epsilon}=2$, i.e., $\epsilon_{i,t}\sim N(0,2)$, $\phi=0.8$ is the autoregressive coefficient, $\boldsymbol{\beta}=(\beta^{(1)},\dots,\beta^{(11)})^{\prime}$ is a $m\times 1$ vector of coefficients and $\mathbf{X}_{i,t-1}=\left(\operatorname{X}_{i,t-1}^{(1)},\dots,\operatorname{X}_{i,t-1}^{(11)}\right)$ is a $1\times m$ vector of $11$ predictors, both continuous and categorical, also containing interaction terms and correlated regressors. We allowed the covariates to vary in time and across units in the dataset by adding, respectively, a random term $\nu_{t}$ (which, as will become clear later, varies for different covariates) and a random term $u_{i}\sim N(1,1)$. The latter term adds variability (and thus, heterogeneity) between the units. In particular, the covariates are generated as follows: $\operatorname{X}_{i,t}^{(1)}=0.1t+u_{i}+\nu_{t}^{(1)},\nu_{t}^{(1)}\sim N(0,1)$, $\operatorname{X}_{i,t}^{(2)}=0.1t+u_{i}+\nu_{t}^{(2)},\nu_{t}^{(2)}\sim N(0,0.2)$, $\operatorname{X}_{i,t}^{(3,4,5)}=u_{i}+\nu_{t}^{(3,4,5)},\nu_{t}^{(3,4,5)}\sim MVN(0,\Sigma)$, $\operatorname{X}_{i,t}^{(6)}=u_{i}-\nu_{t}^{(6)},\nu_{t}^{(6)}\sim N(0,1)$, $\operatorname{X}_{i,t}^{(7)}=\left(0.1t+\nu_{t}^{(1)}\right)^{2}+u_{i}+\nu_{t}^{(7)},\nu_{t}^{(7)}\sim N(0,0.2)$, $\operatorname{X}_{i,t}^{(8)}\in\{0,1\}$, $\operatorname{X}_{i,t}^{(9)}\in\{1,2,3\}$, $\operatorname{X}_{i,t}^{(10)}=\operatorname{X}_{i,t}^{(3)}\cdot\operatorname{X}_{i,t}^{(9)}$, $\operatorname{X}_{i,t}^{(11)}=\operatorname{X}_{i,t}^{(2)}\cdot\operatorname{X}_{i,t}^{(8)}$. [^20] The variance-covariance matrix of the multivariate normal distribution used to generate covariates 3–5 is set to $\Sigma=\begin{bmatrix}1&.5&.7\\ .5&1&.3\\ .7&.3&1\\ \end{bmatrix}$. Notice that, to put the ML algorithms under further stress, the covariance between $\operatorname{X}_{i,t}^{(3)}$ and $\operatorname{X}_{i,t}^{(5)}$ is $0.7$ so the two variables are highly correlated but $\operatorname{X}_{i,t}^{(3)}$ is ten times more important (in terms of coefficient) than $\operatorname{X}_{i,t}^{(5)}$. These choices regarding the data-generating process for the covariates included in our simulation study are driven by the goal of mirroring real-world empirical settings where the methodology might be applied. In addition, to better reflect a typical real-world scenario where only a subset of covariates is relevant, we set certain coefficients to zero: specifically, $\beta^{(1)}$, $\beta^{(8)}$, and $\beta^{(9)}$. The remaining coefficients are generated as follows: $\beta^{(2)}=\beta^{(6)}=\beta^{(10)}=2$, $\beta^{(3)}=\beta^{(7)}=1$, $\beta^{(4)}=2.5$, $\beta^{(5)}=0.1$ and $\beta^{(11)}=1.5$. We also remark that, for computational reasons, in this simulation study we focus on a low-dimensional set of covariates. However, the MLCM is particularly well-suited for data-rich environments and sparse settings with many more covariates. Therefore, the simulation results provided here should be interpreted as conservative evidence regarding the performance of the MLCM. Finally, we assume exogeneity of the predictors in both the pre- and post-intervention periods (third part of Assumption 3). Table 1 below provides an overview of the generated datasets.
@@
At $t_{0}+1$ (corresponding to T = $5$ in Table 1), we included a fictional intervention that increases the outcome for each unit by $2$ standard deviations at time $t_{0}+1$, $1.5$ standard deviations at time $t_{0}+2$, and $1$ standard deviation at time $t_{0}+3$. In other words, the intervention has a decreasing effect over time. We made this choice because adding a unit-specific component in the covariates generates heterogeneity; therefore, the scale of each $\operatorname{Y}_{i}$ varies across the $i$’s. We measured the performance of the MLCM in terms of both the bias of the estimated effect from the true impact and the interval coverage. Also note that, under this setup, the number of pre-intervention periods is $t_{0}=4$ when $T=7$ and $t_{0}=9$ when $T=12$. As estimators, we employ four different ML algorithms - LASSO, Partial Least Squares, random forest, and stochastic gradient boosting.
The results are summarized in Table 2 below and show that, across all post-intervention horizons, the MLCM achieves a very low bias. This holds true for both linear and non-linear model specifications. The bias tends to decrease when the number of pre-intervention time periods increases, as more information is present in the data. Nevertheless, the bias at the shortest time period is still very low, which reinforces our belief that the MLCM can be effectively used for short panels. For instance, under the linear specification and $t_{0}=4$, the relative bias increases from $0.2\%$ measured at $t_{0}+1$ to $0.9\%$ measured at $t_{0}+3$. However, by increasing the number of pre-intervention time points, the bias drops at $0.4\%$ even at $t_{0}+3$. We also observe that the relative bias under a linear model specification is much lower than that under the non-linear specification, which was expected since non-linearities are typically more challenging to detect. In any case, the bias under a non-linear model specification is still low (the maximum bias is $4.9\%$ and is measured in the most challenging scenario, i.e., at the third time horizon when the pre-intervention series is very short, $t_{0}=4$).
@@
[…]
Similar observations can be made for the interval coverage, which was estimated based on $1,000$ block-bootstrap iterations. At the first time horizon after the intervention, the interval coverage is close or equal to the nominal $95\%$ level for both linear and non-linear model specifications. Then, it tends to slightly decrease as we move further away from the intervention, but remains close to the nominal level for both linear and non-linear processes.
In Supplemental Appendix D, we provide many additional results that show that these key findings remain largely unchanged if we consider alternative data-generating processes for both the outcome variable and the covariates, and if we reduce the number of available units. Notably, when $\phi$ is set to $1.2$ (cf. Table D.1), rather than to $0.8$ as in Table 2, i.e., when moving from a stationary to a non-stationary data-generating process with an explosive autoregressive coefficient for the lag of the outcome variable, the MLCM exhibits very similar performance.
@@ 4 Empirical application
The main empirical illustration provided below is an original study on the educational effects of the COVID-19 crisis, thus focusing on a simultaneous treatment affecting all available units. In Supplemental Appendix F, to demonstrate how to leverage the MLCM in settings where there are untreated units, but there may be violations of the no-interference assumption, we revisit the application on minimum wage with staggered treatment adoption in Callaway and Sant’Anna (2021) without using their control group. Both empirical illustrations rely on short panel datasets.
@@ 4.1 Background
In the aftermath of the COVID-19 pandemic, the inequality legacy of this unprecedented crisis has emerged as an issue of great policy relevance. The available evidence on income inequality documents a decrease driven by short-run and temporary government compensation policies, which were mostly targeted at the poorest segments of the population (Stantcheva, 2022). Concerning education, instead, recent micro-level evidence (Agostinelli et al., 2022; Carlana et al., 2023) shows that school closures had a large, persistent, and unequal effect on learning. The educational gaps caused by the school closures may, in turn, permanently affect the lifetime income possibilities of the current generation of students, with vast repercussions on future inequalities (Werner and Woessmann, 2023). Therefore, it is through the human capital channel that the inequality effects of the pandemic may eventually appear in the medium and long run. It follows that to anticipate longer-term consequences on income inequality and territorial disparities, it is necessary to gauge the magnitude of educational losses and their distribution across areas.
To our knowledge, there is no granular evidence on the geography of the education effects of the COVID-19 crisis. This is mainly due to identification challenges caused by the sudden spread of the pandemic across the world, which resulted in the absence of a suitable untreated group. Italy is an important case study, as it ranks among the hardest-hit countries and was the first Western country to impose a strict nationwide lockdown. During the lockdown that began on March 9, 2020, schools across the entire national territory were closed and remained so until the end of the school year. This resulted in Italy ranking among the OECD countries with the highest number of weeks of school closures and distance learning (Battisti and Maggio, 2023). In this setting, we leverage the MLCM with non-staggered adoption, given that the treatment—the COVID-19 shock—simultaneously affected all units. [^21] We refer to the treatment variable as COVID-19 “shock”, which encompasses both the pandemic and the containment policies, as we aim to capture the total effect of the pandemic, i.e., its direct and indirect effects on educational outcomes. In line with previous literature, we consider the restrictive measures, including the lockdown and the school closures, as a manifestation of the pandemic, not as a separate treatment. COVID-related learning losses can originate from many pandemic-induced channels, such as school closures and distance learning, prolonged absence from school due to one or more infections, absence due to parents’ concerns about their children’s health, and many other transmission mechanisms. At the aggregate level, only the overall impact of the pandemic remains discernible.
@@ 4.2 Data and implementation
We employ LLM yearly data covering the period from 2013 to 2020. We cover all Italian LLMs except some of the smallest ones for which the education data are unavailable due to privacy protection, for a total of $579$ LLMs ($95\%$ of Italian LLMs and over $99\%$ of the Italian population). The dependent variable is the standardized math test score of fifth-grade students referred to the school year started in 2020. Following the approach by Carlana et al. (2023), the math score is standardized with respect to its pre-pandemic mean. This way, the detected treatment effects can be interpreted as deviations from the pre-COVID baseline. We focus on younger students for two reasons: i) learning disruptions earlier in life typically have longer-lasting and more severe effects, so younger children may have been more heavily impacted (Stantcheva, 2022); ii) primary schools were excluded from the heterogeneous school closures involved in the tier system of regional restrictions implemented by the Italian government since the onset of the second wave (November 2020), so the treatment is the same for all units (Assumption 1). The policy relevance of the math outcome is illustrated by the fact that the most recent report by the OECD Program for International Student Assessment (PISA), published in December 2023, reported that math scores dropped globally in the last few years, making headlines around the world. [^22] See here for coverage of the news from the New York Times.
The initial pre-treatment information set includes over $150$ variables (see Table LABEL:tabapp:vardef in Supplemental Appendix E for a detailed description of the variables). In this set of covariates, we included the first three lags of all the predictors as covariates and three lags of the outcome variable. This implies that we collapse the original 2013–2020 dataset into a dataset covering the period 2017–2020.
The implementation process for this application is reported in Box 4.2, while the outcomes of the empirical analysis are reported in Section 4.3.
{mdframed}
[ innerlinewidth=0.5pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin = 10pt, innerbottommargin=10pt, frametitle=                      Box 4.2: Estimating the local impact of the COVID-19 shock on education, frametitlerule=true, frametitlerulewidth=1pt]
Preliminary
  1. 0.
    Data splitting. We split the full 2017-2020 dataset into two subsets according to the treatment date: 2017-2019 to be used in the Design Stage; 2020 to be used in the Analysis Stage.
    Design Stage
  2. 1.
    Algorithm selection. We select four supervised ML algorithms (the same set used for the simulation of Section 3.3): 1) LASSO; 2) Partial Least Squares; 3) stochastic gradient boosting; 4) random forest. As we are agnostic about the functional form of the underlying data-generating process, we opt for a mix of non-linear and linear models.
  3. 2.
    Principled input selection. We build an initial LLM dataset with over $150$ predictors on the basis of literature insights. From this dataset, we then keep only the most important predictors selected by a preliminary random forest run on the pre-treatment data (see step below).
  4. 3.
    Panel cross-validation. For each algorithm, we tune hyperparameters via panel CV, involving iterative estimation on two different training-testing pairs of pre-COVID datasets: i) 2017 training; 2018 testing; ii) 2017-2018 training, 2019 testing.
  5. 4.
    Performance assessment. We assess average performance metrics for the four algorithms by comparing average forecasted vs. actual outcomes on the 2018–2019 held-out test data. We then compare the performance of the different MLCM versions.
  6. 5.
    Diagnostic and placebo tests. We first check the average distribution of errors with the best-performing model for the 2018-2019 testing sets and then show the map of the unit-level placebo temporal average treatment effects in the pre-COVID period.
    Analysis Stage
  7. 6.
    Final model selection. On the basis of the comparative performance assessment, we pick the best-performing algorithm (random forest), with its best configuration, and re-train it on the full 2017–2019 sample using the hyperparameters cross-validated in the Design stage.
  8. 7.
    Counterfactual forecasting. We apply the final model estimated in Step 6 and forecast, for each LLM $i$, the counterfactual outcome $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0)$.
  9. 8.
    Estimation of treatment effects. For each LLM $i$, we estimate the individual treatment effect by taking the difference between the observed post-COVID outcome $\operatorname{Y}_{i,t_{0}+1}$ and the ML-generated potential outcome $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0)$.
  10. 9.
    Treatment effect heterogeneity. We estimate data-driven CATEs via a regression tree analysis with the individual treatment effects as the outcome and a host of potentially relevant predictors associated with the heterogeneity of the education effects.
  11. 10.
    Inference. We estimate standard errors for the ATE and CATEs via block-bootstrapping by performing $1,000$ bootstrap replications of Steps 6 to 9.
@@
We use a data-driven approach to restrict the information set. More specifically, we follow Athey and Wager (2019) and Basu et al. (2018) and apply a pilot random forest on the pre-treatment data to pick up a subset of the most important predictors according to the importance ranking produced by the forest. [^23] For this preliminary operation, we use default hyperparameter settings that typically perform well with random forests (Athey and Wager, 2019). See Figure E.1 for the variable importance ranking of the preliminary forest. In order to select the precise number of relevant predictors in a data-driven manner, we include this number as an additional parameter in the subsequent panel CV routine which we employ to tune the hyperparameters of all the selected ML algorithms (LASSO, Partial Least Squares, stochastic gradient boosting, and random forest ). [^24] For all algorithms, we tune the subset of most predictive variables (10, 20 or 30) to be included in the analysis according to the preliminary forest. For LASSO, we tune the penalty parameter (the values from 0.1 to 0.9 with an increase of 0.1 in each run); for PLS, we tune the number of components (all possible values from 1 to the number of variables). For boosting, we tune the number of trees (1,000 or 2,000), the maximum depth of each tree (1 or 2), the minimum number of observations in terminal nodes (5 or 10) and the learning rate (0.001, 0.002 or 0.005); for random forest, we tune the number of variables randomly sampled as candidates at each split (1/2, 1/3 or 1/4 of the number of variables), and use a fixed number of 1,000 trees. All candidate hyperparameter values are tested on each testing set.
@@
Note that in this application, due to the ubiquitous nature of this shock, we do not invoke the third part of Assumption 3 (exogeneity of post-treatment covariates) and only include pre-treatment values of time-varying variables and time-invariant predictors in the information set. Assumption 3 is satisfied because the pandemic was completely unanticipated. Regarding Assumption 4, given the unprecedented disruptions brought about by the pandemic and based on institutional and subject matter knowledge, we deem it plausible that any divergence from the forecasted counterfactual in educational outcomes can be attributed exclusively to the COVID-19 shock, and not to other concomitant shocks or policies. [^25] There were no educational reforms at the primary school level that could have influenced the standardized math test scores of fifth-grade students in the period under scrutiny. Finally, Assumption 5 is not needed for identification, as we only estimate one-step-ahead causal effects here.